OpenStudy (anonymous):

f is a differentiable function on the interval [0, 1] and g(x) = f(4x). The table below gives values of f ′(x). What is the value of g ′(0.1)?

OpenStudy (anonymous):

https://gyazo.com/b435133c157b25d1c60f85c59cdeba34

OpenStudy (anonymous):

Just multiplied by 4. Would that be right?

OpenStudy (anonymous):

@zepdrix it's destiny telling you to help me ;P

zepdrix (zepdrix):

hehe

zepdrix (zepdrix):

\[\large\rm g(x)=f(4x)\]Take derivative, apply chain rule on the right side of the equation.

zepdrix (zepdrix):

Understand that step? :d

OpenStudy (anonymous):

Chain rule on th eright side?

zepdrix (zepdrix):

\[\large\rm \frac{d}{dx}f(4x)\quad=f'(4x)\cdot(4x)'\]Yes, chain rule. Multiply by the derivative of the inner function.

zepdrix (zepdrix):

\[\large\rm g(x)=f(4x)\]Differentiate with respect to x,\[\large\rm \frac{d}{dx}g(x)=\frac{d}{dx}f(4x)\]Leading to\[\large\rm g'(x)=f'(4x)\cdot(4x)'\]this\[\large\rm g'(x)=f'(4x)\cdot(4)\]

zepdrix (zepdrix):

Confused?

OpenStudy (anonymous):

I got it just confused on how that'll help me.

zepdrix (zepdrix):

Well we're looking for g'(0.1). Let's plug that into this relationship.\[\large\rm g'(0.1)=4\cdot f'(4\cdot0.1)\]\[\large\rm g'(0.1)=4\cdot f'(0.4)\]

zepdrix (zepdrix):

And then use the table to finish the problem :o

OpenStudy (anonymous):

Ohhhhhhhhhhhhhh Got it -16

zepdrix (zepdrix):

Yay!

OpenStudy (anonymous):

8/8 m8 that was gr8 thx

zepdrix (zepdrix):

lol

OpenStudy (anonymous):

Good job m8